# StableMaterials: Enhancing Diversity in Material Generation via Semi-Supervised Learning

### v2

@@ 3.4. Fast High-Resolution Generation
To enhance generation efficiency, we finetune a Latent Consistency Model [Luo et al., 2023] (LCM). This model streamlines the denoising process by implementing a one-stage guided distillation of an augmented probability flow ordinary differential equation (PF-ODE). The distillation leverages a consistency function $f_{\theta}:(z_{t},c,t)\mapsto z_{0}$, which directly predicts the PF-ODE solution at $t=0$. Compared to two-stage methods [Meng et al., 2023], LCM integrates the guidance scale $\omega$ directly into the augmented PF-ODE. Additionally, LCM adopts a skipping step strategy, ensuring consistency not between consecutive steps, but between steps separated by $k$, thus $t_{n+k}\rightarrow t_{n}$, circumventing alignment issues in traditional models. The use of consistency models enables few-steps generation, resulting in fast material synthesis as detailed in Sec. 4.2.
@@
Recent material generation methods achieve tileability via noise rolling [Vecchio et al., 2023], leveraging the iterative nature of diffusion models. However, this approach becomes ineffective with fewer generation steps, being unable to fully remove seams. Circular padding [Zhou et al., 2022], on the other hand, despite being independent of diffusion steps, is not suitable for patched generation as it causes self-tileability in patches.
To address these issues, we propose shifting tensor rolling from the diffusion step to the U-Net architecture by directly rolling the feature maps. This method, named features rolling, integrates rolling within U-Net blocks, reducing the need for multiple diffusion steps. By randomly translating and then reversing feature maps within each convolution and attention layer in the U-Net, this approach preserves edge continuity and minimizes visual artifacts typically associated with fewer diffusion steps. To better preserve structure in highly structured materials, we enable features rolling after the first diffusion step. We compare features rolling with other approaches to tileability in the Supplemental Materials.
@@
Features rolling in combination with patch diffusion allows efficient high-resolution generation; however, it results in inconsistent generations between patches, due to the loss of cross-patch dependencies.
To overcome this limitation, we use a two-stage pipeline, similar to the one introduced in SDXL [Podell et al., 2023], where the output of the base model is further refined by a high-resolution expert using SDEdit [Meng et al., 2021]. This approach allows us to use patched diffusion in the refinement process while keeping the overall content consistent with the original generation. In particular, we first generate at the model’s base 512x512 resolution, and then refine it at the desired resolution. We apply the SDEdit refinement with a strength of $0.5$, which we find to be the best compromise between new high-frequency details cretion and consistency with the previous step generation. To better learn the high-frequency features of the materials, the refiner model is trained on 512x512 crops extracted from the 4K materials at their full resolution without downsampling. We show the quality improvement of this approach in the ablation study, Sec. 4.4.3, and provide additional samples in Supplemental Materials.
@@ 4.2. Technical details
The compression model is trained with mini-batch gradient descent, using the Adam optimizer [Kingma and Ba, 2014] and a batch size of $8$. We train the model for $1,000,000$ iterations with a learning rate of $4.5\cdot 10^{-4}$ and enable the $\mathcal{L}_{\text{adv}}$ after $300,000$ iterations as in Esser et al. [2021]. We first train a multi-encoder VAE [Vecchio et al., 2024]; then we fine-tune a single-encoder model to compress the concatenated maps into the same disentangled latent space, reducing the total of parameters from 271M to 101M, while keeping similar reconstruction performance (Tab. 1). We fine-tune both the single encoder model and the texture encoder for $100,000$ steps while keeping the decoder frozen.
@@
Latent Consistency model
Inference
We assess execution speed and memory usage. We generate with 4 denoising steps, followed by 2 refinement steps, using the LCM sampler with a fixed seed and processing up to 8 patches in parallel. Both base and refiner models run at half-precision. Generation takes 0.6s at $512\times 512$, 1.5s at $1024\times 1024$ and 6.5GB of VRAM, 4.9s at $2048\times 2048$ and 7.4GB VRAM, and 18.6s at $4096\times 4096$ and 12GB of VRAM. For comparison, the LDM with a DDIM sampler and 50 generation steps, followed by 25 refinement steps, takes 20.6s at $2048\times 2048$, and 65.4s at $4096\times 4096$.
@@ 4.3. Results and comparison
For all results, we show generated material maps, as well as the rendering under ambient lighting. We include a wider variety results in the Supplemental Material, as well as material editing samples and a CLIP-based Nearest-Neighbour search in the training database.
@@
We evaluate our model generation capabilities for both image (Fig. 4 and Fig. 12) and text (Fig. 5 and Fig. 13) conditioning. The results show the ability of StableMaterials to closely follow the input prompt, producing diverse realistic materials. For each set of results, we show two in-domain generations, that is materials categories that are present in our annotated training dataset, and two out-domain samples, materials from the unannotated data, confirming the ability of StableMaterials to generate new high-quality materials from both data distributions.
Comparison
We compare StableMaterials against MatFuse [Vecchio et al., 2024] trained on MatSynth [Vecchio and Deschaintre, 2024], MatGen [^1] Generation samples for MatGen were kindly provided by the authors.  [Vecchio et al., 2023], Material Palette [Lopes et al., 2024], and the text-to-texture-to-material pipeline in Adobe Substance 3D Sampler [Adobe, 2024]. Figures 6 and 7 show the generation results for image and text conditioning, respectively. It should be noted that MatFuse, Material Palette, and StableMaterials are trained on public datasets, while both MatGen and Sampler use private data. Furthermore, MatFuse, MatGen, and StableMaterials directly produce PBR maps, while MaterialPalette and Sampler employ a two-step process: first, generating a texture, then using an SVBRDF estimator to predict the material maps. This approach benefits from leveraging large image models, but suffers from biases inherent in SVBRDF predictions, which guess material properties from light-surface interactions, that are further exarcebated by the non-material-specific training sometimes generating natural images (with perspective) instead of surfaces. Moreover, MaterialPalette fine-tunes a LoRA for each prompt, introducing a time and computation overhead. In contrast, StableMaterials directly creates material maps by having learned material features, leading to more precise and unbiased results. MatFuse is constrained by its $256\times 256$ resolution, resulting in blurry, undetailed generations, and struggles with complex materials. MatGen, on the other hand, generally synthesize high-quality materials, but presents oversharpening artifacts and struggles in following more complex promps. We only show the generation renderings in Figures 6 and 7, with the PBR maps and additional samples included in the Supplemental Materials.
@@ 4.4.2. Training strategies
We evaluate the semi-supervised training on model’s generation capabilities in Fig. 8. The base model (w/o semi-supervision) generates high-quality in-domain material but, as expected, struggles with out-domain materials absent from the original training data. The semi-supervised approach, in contrast, enhances diversity and overall quality, allowing the generation of materials beyond the initial training scope.
@@
W/o Semi-supervision W/ Semi-supervision W/o Semi-supervision W/ Semi-supervision
@@
Base Refined Base Refined
@@
Figure 9. Refinement ablation results. Generation crops: the use of a diffusion refiner significantly enhances generation quality and sharpness.
@@ 4.4.3. High-resolution generation
We evaluate our method with and without the refiner in Fig. 9. The generation looks sharper, less noisy, and more detailed. We demonstrate the effectiveness of our two-stage approach to high-resolution generation in Fig. 10, comparing it to patched diffusion, complemented by feature rolling. Patched diffusion leads to increased scale (pattern replication) rather than enhanced resolution (more pixel density) and causes inconsistencies like scale variations between patches. In contrast, our approach generates consistent and better quality high-res materials by refining the generation at the model’s native resolution of $512\times 512$ pixels via a second diffusion step at the target resolution.
Patch diffusion Two-stages pipeline
Refer to caption Refer to caption
‘Tortoise shell.’
Figure 10. Ablation of the approaches to high-resolution. (1) Patched diffusion produces more scale rather than resolution and presents inconsistency in scale. (2) Our two-stage approach generates at the model’s native resolution, and refines the material.
@@ 5. Limitations and Future Work
StableMaterials presents some limitations as shown in Fig. 11. First, it struggles to handle natural prompts describing spatial relations and is unable to accurately represent complex concepts or figures. Introducing more variety in training prompts could help mitigate the problem. Additionally, it can occasionally generate incorrect reflectance properties (e.g.: material misclassified as metal) for material classes that are only present in the unannotated dataset. The use of text prompts describing surface properties, at training time, could mitigate the problem.
@@ 6. Conclusion
We present StableMaterials, a novel diffusion model for fast generation of tileable high-resolution materials. Our model leverages semi-supervised training and knowledge distillation from large-scale pre-trained models to overcome the limitations in training data and improve material realism and variety. This approach significantly reduces reliance on annotated data, thereby enriching the diversity and realism of generated PBR materials.
In conclusion, StableMaterials presents a robust solution to material generation challenges and we believe it can serve as a blueprint for future research to harness the potential of large-scale pretrained models and unsupervised data in computer graphics and beyond.

### v3

@@ 3.4. Fast High-Resolution Generation
To improve generation speed, we fine-tune a Latent Consistency Model (LCM) [Luo et al., 2023]. LCM performs a one-stage guided distillation of an augmented Probability Flow ODE (PF-ODE) and directly predicts the solution at $t=0$ through a consistency function $f_{\theta}(z_{t},c,t)\mapsto z_{0}$. Unlike two-stage methods [Meng et al., 2023], LCM integrates the guidance scale $\omega$ directly into the PF-ODE and uses a skip-step strategy, ensuring consistency between time steps $t_{n+k}$ and $t_{n}$. This design avoids alignment issues present in typical two-stage approaches. As a result, LCM enables generation in only a few steps, resulting in faster material synthesis.
@@
Noise rolling [Vecchio et al., 2024a] has been used to achieve tileability through iterative diffusion but becomes less effective with fewer steps. We address this limitation proposing the features rolling, which shifts rolling from the noisy inputs to the U-Net features. Within each convolution and attention layer, we randomly roll and then reverse the feature maps, thus preserving edge continuity while requiring fewer diffusion steps. For highly structured materials, we enable features rolling only after the first diffusion step to retain their global layout. We compare features rolling with other tileability methods in the Supplemental Materials.
@@
Combining features rolling with patch-based diffusion provides efficient, high-resolution synthesis, but patch-wise generation alone can introduce inconsistencies across tiles. We address this with a two-stage pipeline, similar to SDXL [Podell et al., 2023], using SDEdit [Meng et al., 2021] for refinement. Specifically, we first generate at $512\times 512$ resolution, then refine the output to the desired resolution with a strength of $0.5$, balancing new high-frequency details against global consistency. The refiner model is trained on full-resolution $512\times 512$ crops from 4K materials (no downsampling), ensuring it captures fine surface details. We demonstrate the effectiveness of this approach in Sec. 4.4.3, with additional samples provided in the Supplemental Materials.
@@ 4.2. Technical details
The compression model is trained with mini-batch gradient descent, using the Adam optimizer [Kingma and Ba, 2014] and a batch size of $8$. We train the model for $1,000,000$ iterations with a learning rate of $4.5\cdot 10^{-4}$ and enable the $\mathcal{L}_{\text{adv}}$ after $300,000$ iterations as in Esser et al. [2021]. We first train a multi-encoder VAE [Vecchio et al., 2024b]; then we fine-tune a single-encoder model to compress the concatenated maps into the same disentangled latent space, reducing the total of parameters from 271M to 101M, while keeping similar reconstruction performance (Tab. 2). We fine-tune both the single encoder model and the texture encoder for $100,000$ steps while keeping the decoder frozen.
@@
Latent Consistency model
Inference
We assess execution speed and memory usage. We generate with 4 denoising steps, followed by 2 refinement steps, using the LCM sampler with a fixed seed and processing up to 8 patches in parallel at half-precision. Generation takes 0.6s at $512\times 512$, 1.5s at $1024\times 1024$ and 6.5GB of VRAM, 4.9s at $2048\times 2048$ and 7.4GB VRAM, and 18.6s at $4096\times 4096$ and 12GB of VRAM. In contrast, an LDM with a DDIM sampler (50 steps) plus 25 refinement steps requires 20.6s at $2048\times 2048$, and 65.4s at $4096\times 4096$.
@@ 4.3. Results and comparison
All results show both generated material maps and ambient-lit renderings. For evaluation, we carefully selected test prompts to include both in-domain concepts (present in training categories) and novel out-domain concepts, ensuring no direct overlap with training prompts. Additional samples, material editing results, and a CLIP-based nearest-neighbor search in our training database are included in the Supplemental Materials.
@@
We show the generative capabilities of our model, for both image (Fig. 4) and text (Fig. 5) conditioning. We include both in-domain samples (categories found in the annotated dataset) and out-domain samples (from unannotated data). In all cases, StableMaterials produces realistic results that closely follow the prompts.
Qualitative comparison
We compare StableMaterials against MatFuse [Vecchio et al., 2024b], MatGen [Vecchio et al., 2024a], Material Palette [Lopes et al., 2024], and Adobe Substance 3D Sampler [Adobe, 2024], as shown in Figures 6 and 7 for image and text prompting, respectively. MatFuse, Material Palette, and StableMaterials each use public datasets, while MatGen and Sampler rely on private data. MatFuse is limited to a $256\times 256$ resolution, often generating blurry or simplistic outputs and struggling with complex textures. Material Palette and Sampler take a two-step approach (texture generation followed by SVBRDF estimation).This approach benefits from leveraging large image models, but suffers from biases inherent in SVBRDF predictions, which guess material properties from light-surface interactions, that are further exacerbated by the nonmaterial-specific training sometimes generating natural images (with perspective) instead of surfaces. Moreover, MaterialPalette fine-tunes a LoRA for each prompt, introducing a time and computation overhead. MatGen, on the other hand, produces high-quality materials, but presents over-sharpening artifacts and struggles in following more complex promps. In contrast, StableMaterials directly outputs PBR maps using a tailored latent representation and a two-stage generation process, mitigating resolution constraints and reducing artifacts. We only show the generation renderings in Figures 6 and 7, with the PBR maps and additional samples included in the Supplemental Materials.
Quantitative Comparison
Material quality is difficult to assess with standard image metrics (FID [Heusel et al., 2017], IS [Salimans et al., 2016]) due to the different data distribution from that of natural images. Instead, we leverage CLIP-based metrics (CLIP Score [Hessel et al., 2021] and CLIP-IQA [Wang et al., 2023]), which assess both semantic alignment with prompts and perceptual quality of outputs. Table 1 shows StableMaterials either outperforms or matches state-of-the-art methods—some trained on larger datasets—and significantly improves over MatFuse. Comparison is carried out on 80 text-conditioned generation.
Table 1. Quantitative comparison. We compare the generation quality of StableMaterials to MatFuse, MatGen and Substance 3D Sampler. The CLIP-IQA score is computed using the “high-quality/low-quality” contrastive pair.
MatFuse MatGen Substance 3D Sampler Stable Materials
CLIP Score $\uparrow$ 26.2 28.8 24.9 29.6
CLIP-IQA $\uparrow$ 0.52 0.66 0.71 0.70
@@ 4.4.2. Training strategies
We evaluate the effect of the semi-supervised training on model’s generation capabilities in Fig. 8. Our semi-supervised training improves generation diversity, especially on out-of-domain materials that the purely supervised baseline cannot produce reliably.
@@
W/o distillation W/ distillation W/o distillation W/ distillation
@@
Base Refined Patch diffusion Two-stages pipeline
@@
(a) Refinement ablation.
(b) High-resolution ablation.
Figure 9. Ablation studies comparing (a) the effect of refinement on quality and (b) different approaches to achieving high-resolution. Results show that the use of a diffusion refiner significantly enhances generation quality and sharpness, and two-stage approach generates at the model’s native resolution avoids the inconsistency in scale of the patched diffusion.
@@ 4.4.3. High-Resolution Generation
Our two-stage approach to high-resolution sharpens and improves details (Fig.9(a)). Additionally, using only patched diffusion can introduce scale and consistency artifacts across patches, whereas our two-stage approach (base generation + refinement) produces more coherent large-format outputs by consolidating patches at $512\times 512$ before upscaling (Fig.9(b)).
@@ 5. Limitations and Future Work
StableMaterials presents some limitations as shown in Fig. 10. First, it struggles with natural prompts describing spatial relations and is unable to accurately represent complex concepts or figures. Introducing more variety in training prompts could help mitigate the problem. Additionally, it can occasionally generate incorrect reflectance properties (e.g.: material misclassified as metal) for material classes that are only present in the unannotated dataset. The use of text prompts describing surface properties, at training time, could mitigate the problem. Finally, despite being able to represent a wide variety of materials outside the annotated dataset, it is still limited to the classes represented in the unannotated data.
@@ 6. Conclusion
We introduce StableMaterials, a novel diffusion-based model for fast, tileable, high-resolution material generation. By integrating semi-supervised training and knowledge distillation from large-scale pretrained models, StableMaterials overcomes the lack of annotated data and delivers greater realism and variety in PBR materials. Moreover, feature-rolling enables tileable generation in few-steps settings, making our approach practical for real-world applications.
We believe that StableMaterials can serve as a blueprint for future research, demonstrating how domain-specific generators can effectively leverage large-scale pretrained models and unsupervised data to improve diversity while maintaining physical plausibility and generation quality.